OpenStudy (anonymous):

What is the value of x? 1/4x - 8 + 2/4x = 19

OpenStudy (anonymous):

@ConDawg Can you help me?

OpenStudy (anonymous):

\[\frac{ 1 }{ 4}x- 8 +\frac{ 2 }{ 4 }x = 19\]

OpenStudy (anonymous):

Step One: Add the 1/4x and the 2/4x. \[\frac{ 1 }{ 4 }x + \frac{ 2 }{ 4 }x = \frac{ 3 }{ 4 }x\] Now we have \[\frac{ 3 }{ 4 }x -8 = 19\]

OpenStudy (anonymous):

Step Two: Move the -8 over to the other side. \[\frac{ 3 }{ 4 }x = 27\]

OpenStudy (anonymous):

Step Three: multiply both sides by 4 \[3x = 108\]

OpenStudy (anonymous):

Last Step: Divide both sides by 3. \[x = 36\] So, that's our answer. x = 36

OpenStudy (anonymous):

I have to study for an exam. So that's all I can help with for now. Good luck.